You know that the assets of a firm SKIP are today worth 100mil. You reasonably feel that in a year they will be either worth 110mil or 90mil. You also know that a treasury bill maturing in one year is offering today a yield of 5%. The firm has a zero-coupon bond that matures in one year and has a face value of 100mil. What should be the value of this corporate bond today? What should be its yield to maturity? What should be the value of the equity of the firm?
Answer #1
  1. Value of bond= Face vale / (1+ yield )^n

Here face value = 100 mil

Yield= 5%

N= 1

So the Value = 100/(1+0.05)

= $93.24

  1. YTM of zero coupon bond = face value/ current price)^(1/ time period) -1

So , YTM will be = (100/93.24)1-1

                = 1.0725-1

                = 0.0725 or 7.25%

So the YTM is 7.25%

  1. Even if the bonds are issued a discount, the bonds are entered at its face value in the books so here the value of bond in the balance sheet = $100 mil and it is given that the assets of SKIP are worth $100mil. This means that there is no equity capital in this firm , only Debt funding.